How fast will a 0.15kg hockey puck accelerate if it is hit with 1.2 N of force?
1
Expert's answer
2020-11-25T10:52:14-0500

By the definition of the Newton's Second Law of Motion, we have:


"F=ma,""a=\dfrac{F}{m}=\dfrac{1.2\ N}{0.15\ kg}=8\ \dfrac{m}{s^2}."

Answer:

"a=8\ \dfrac{m}{s^2}."